OpenStudy (anonymous):

Simplify with the the check mark thing again the number 0.16. a) 4 b) .08 c) 0.04 d) 0.4 ill give to the best answer plz help me. @LastDayWork

OpenStudy (lastdaywork):

0.16 = 16/100 Now; 4*4 = 16 & 10*10 = 100 Can you now answer the question ??

OpenStudy (anonymous):

0.04 or 0.4 right

OpenStudy (lastdaywork):

Answer in the form of a fraction.

OpenStudy (anonymous):

ohh its 4

OpenStudy (lastdaywork):

Actually the answer is 4/10 = 0.4

OpenStudy (anonymous):

oh ok i get it now thank you

OpenStudy (lastdaywork):

I'll advice you to solve in terms of fraction..this way you never make a mistake with decimal point :)

OpenStudy (anonymous):

ok thank you

OpenStudy (lastdaywork):

:)